OpenStudy (xsamx3):

Solve using matrices. 5x – 6y = 50 -x + 5y = -48

OpenStudy (xsamx3):

Need help with this! I need someone to show me step by step, then ill try the others.

OpenStudy (unklerhaukus):

\[5x-6y = 50\\ -x +5y = -48\] \[\begin{bmatrix} 5&-6\\ -1&5\\ \end{bmatrix}\begin{bmatrix} x\\y \end{bmatrix}=\begin{bmatrix} 50\\-48 \end{bmatrix}\] \[\begin{bmatrix} 5&-6&|&50\\ -1&5&|&-48\\ \end{bmatrix}\]

OpenStudy (unklerhaukus):

Now, what row operation will get a 1, into the top left element of the matrix?

OpenStudy (wolf1728):

5x – 6y = 50 -x + 5y = -48 denominator = 5 -6 -1 5 denom = 25 -6 = 19 x = 50 -6 -48 5 x = 250 -288 / denom x = -38 / 19 x = -2

OpenStudy (unklerhaukus):

\[R_1\to R_1+4R_2\\ \begin{bmatrix} 1&14&|&-142\\ -1&5&|&-48\\ \end{bmatrix}\] \[R_2\to R_2+R_1\\ \begin{bmatrix} 1&14&|&-142\\ 0&19&|&-190\\ \end{bmatrix}\] \[R_2\to R_2/19\\ \begin{bmatrix} 1&14&|&-142\\ 0&1&|&-10\\ \end{bmatrix}\]

OpenStudy (unklerhaukus):

\[R_1\to R_1-14R_2\\ \begin{bmatrix} 1&0&|&-2\\ 0&1&|&-10\\ \end{bmatrix}\] \[\begin{bmatrix} 1&0\\ 0&1\\ \end{bmatrix}\begin{bmatrix} x\\y \end{bmatrix}=\begin{bmatrix} -2\\10 \end{bmatrix}\]